# Observer–Actor: Active Vision Imitation Learning with Sparse-View Gaussian Splatting

### v1

@@ I Introduction
Current imitation learning methods for robotic manipulation primarily rely on static cameras [chi2023diffusion, doi:10.1126/scirobotics.adv7594], egocentric wrist-mounted cameras [chi2024universal, di2024dinobot], or a combination of both [zhao2023learning, wang2025one]. Static cameras are typically placed in a task-agnostic manner, and once a policy is trained, it often must be deployed in the same setup. Wrist-mounted cameras offer greater flexibility but suffer from limited global awareness and a restricted field of view. Combining both viewpoints can partially mitigate these issues but often introduces redundant observations that may distract the model—especially in low-data regimes [chuang2024active, xiong2025vision]. These limitations are further amplified by occlusions, highlighting the need for active vision, where the camera can move dynamically to provide the best viewpoint for the task.
Realizing active vision requires both movable camera hardware and sophisticated active vision strategies. Recent studies [chuang2024active, xiong2025vision, pmlr-v270-cheng25b] achieve this by developing a dedicated active-vision arm for perception, with its policy learned through teleoperation and behavior cloning. However, this approach has two significant limitations. First, the active vision arm is restricted to a fixed observer role, limiting its reachable viewpoints and preventing its use as regular manipulator. Second, the active vision strategy demands extensive human demonstrations, which, although collectable via a VR-based teleoperation system, imposes additional burden on operators.
In this paper, we present ObAct (short for Observer–Actor), a novel framework for active vision in imitation learning, where an observer robot computes optimal visual observations that guide the actor robot’s actions. As illustrated in Figure LABEL:fig:intro with a mug-handle pickup task, ObAct positions the observer arm at test-time optimal viewpoints, providing visual observations for the actor arm during manipulation. Unlike prior approaches, ObAct dynamically assigns observer and actor roles based on the scene configuration, without requiring a separately trained active vision strategy for each arm. To enable this, we construct a test-time 3D Gaussian Splatting (3DGS) representation [kerbl20233d] from sparse-view images and optimize the camera pose to obtain observations that are similar to the demonstration observations, while minimizing occlusions. We further extend existing imitation learning methods—trajectory transfer [doi:10.1126/scirobotics.adv7594] and behavior cloning [zhao2023learning]—to leverage this view-conditioned framework. Experiments show that ObAct substantially improves success rates over static-camera baselines: trajectory transfer increases by 145% without occlusion and 233% with occlusion, while behavior cloning increases by 75% and 143%, respectively.
@@ II Related Work
Active Vision for Robotic Manipulation. Active vision (AV) [aloimonos1988active] refers to actively moving the camera to maximize perceptual quality for downstream tasks, and has been widely applied in robotics for object recognition [7139782], pose estimation [7139782, an2024rgbmanip], and 3D reconstruction [7487527, lee2022uncertainty]. It has also been extensively studied in robotic grasping, where unobserved views are sampled and evaluated using depth images based on information gain [breyer2022closed], grasp affordances [pmlr-v229-zhang23i], graspness inconsistency [NEURIPS2024_4364fef0], and velocity fields [shi2025viso], yielding improved performance in occluded and cluttered environments. In contrast, our method renders and evaluates unobserved viewpoints directly in the RGB domain, the prevalent modality in modern imitation learning systems [chi2023diffusion, zhao2023learning], and extends beyond grasping to support a broader range of manipulation tasks. Our setup is most closely related to [wang2025observe], where an agent first observes the environment and then interacts with it by jointly training a next-best-view camera policy and a next-best-pose gripper policy using few-shot reinforcement learning. However, their approach assumes a freely moving camera and is validated only in simulation, whereas we demonstrate a real-world dual-arm system that achieves active vision without requiring a separate policy. In concurrent work, [yang2025mobipi] also avoid training by leveraging 3DGS to build a scene model from dense scans and applying Bayesian optimization to determine an initial robot pose for task execution on a mobile platform. However, our focus is on sparse-view inputs in a dual-arm setting. [kerrj2025eyerobot] capture demonstrations with a 360° camera and jointly train an eyeball gaze controller and arm policies using a BC–RL loop, but their 2-DoF gaze mechanism is less capable of handling occlusions compared to our 6-DoF observer arm.
@@
View-conditioned Imitation Learning. For manipulation systems to generalize, visuomotor policies must remain robust to discrepancies in observational viewpoints—particularly unseen camera poses—which is especially critical in active vision settings where camera positions are dynamic rather than static. Recent works address this by training end-to-end behavior cloning policies that jointly control both the active camera and the manipulation arms using human demonstrations [xiong2025vision, pmlr-v270-cheng25b, chuang2024active]. While effective, these approaches increase data requirements and model complexity. In contrast, our approach trains policies over a compact distribution of viewpoints and explicitly optimizes the test-time viewpoint relative to demonstrations. Another class of methods builds explicit 3D scene representations, such as point clouds or voxels [shridhar2023perceiver, ze20243d, gervet2023act3d], or renders canonical 2D views from these structures [goyal2023rvt]. These approaches improve robustness to viewpoint variation but require precise extrinsic calibration and incur high computational costs. Data augmentation offers a complementary direction: [tian2024view] generate multi-view datasets from single-view demonstrations using a pre-trained zero-shot novel-view synthesis model [sargent2023zeronvs], while [yang2025novel, ren2025learning] edit reconstructed 3D scenes built from dense scans to create synthetic training views. Beyond end-to-end training, some works transfer a single demonstration trajectory at test time by estimating relative object pose changes between demonstration and execution views, either via pose estimation [doi:10.1126/scirobotics.adv7594] or visual servoing [wang2025one]. However, such methods remain highly sensitive to occlusions. In this paper, we extend behavior cloning and trajectory transfer within our observer–actor framework, demonstrating consistent performance gains in both non-occluded and occluded scenarios.
@@ III-B Sparse-view Gaussian Splatting
Role Assignment. Roles are assigned to each arm based on how closely their captured views match $v^{*}_{\mathrm{demo}}$. Instead of performing explicit relative pose estimation, we use the robust dense feature matcher RoMa [edstedt2024roma], using the number of confident matches on the segmented object as a proxy. The arm with more aggregated matches is designated as the observer, indicating it is closer to the demonstration view, while the other arm becomes the actor.
Three-View GS. We reconstruct the scene using the three images captured by the observer, via InstantSplat [fan2024instantsplat], a sparse-view method that leverages geometric priors. It employs Mast3R [leroy2024grounding], a large-scale 3D geometric model, to estimate camera poses from RGB images and perform joint optimization. These optimized poses are more accurate than those from our low-cost robot arms, resulting in higher-quality reconstructions. We use only three images to train the 3DGS model, as the additional views provide little extra information for rendering $v^{*}_{\mathrm{test}}$, and training is considerably faster. We demonstrate that three views offer the best trade-off between performance and computation time (Section IV-C).
@@ III-C View Optimization
Candidate View Sampling. To initialize the optimization, we sample a hemisphere of candidate viewpoints around the task object’s center, following [breyer2022closed]. The object center is estimated from its 3D bounding box, obtained by lifting 2D masks from GroundedSAM [ren2024grounded] into 3D using aligned depth maps. After filtering out kinematically infeasible poses, we obtain a candidate set $\mathcal{V}_{\mathrm{candidate}}\subseteq\mathcal{V}$. Owing to the efficiency of 3DGS, all candidate views can be rendered at $\sim$250 FPS.
Optimal View Initialization. After rendering a set of virtual RGB images, we select the candidate view with the highest score as the initialization. The scoring function, identical to that used for role assignment, is based on the number of confidence-weighted feature matches between images from $\mathcal{V}_{\mathrm{candidate}}$ and $v^{*}_{\mathrm{demo}}$. This criterion implicitly accounts for occlusion, as occluded regions naturally produce fewer and less reliable matches. In practice, we found that this matching-based strategy outperforms an alternative based on DINOV2 [oquab2024dinov2] cosine distance for measuring view similarity. Even when combined with object segmentation, DINOV2 struggles to discriminate between viewpoints—an observation consistent with [NEURIPS2024_23866f14]. To reduce computational cost, we employ Tiny RoMa [edstedt2024roma], a lightweight variant of RoMa with faster runtime.
Differentiable Rendering. After obtaining an initialization, we refine the viewpoint $v$ using differentiable rendering while explicitly accounting for gripper-induced occlusions. This is particularly important in our setting, where the observer arm’s gripper consistently remains in the field of view and may block the object, as shown in Figure 3. To handle this, we use SAM2 [ravisam] for segmentation of the rendered image, guided by the demonstration image’s mask as a prompt. This is effective because the initialized and demonstration viewpoints are already closely aligned. To mitigate errors in estimating the object center, we first re-center the object’s mask in the image plane. We then refine $v$ by optimizing a loss that aligns the image rendered from the 3DGS model (denoted $\mathcal{G}$) with the segmented demonstration image $I^{*}_{\mathrm{demo}}$, while penalizing occlusions. Let $I(v;\mathcal{G})$ denotes the rendered image. The loss $\mathcal{L}$ as a particular view $v$ is defined as:
@@
where ${\phi}(\cdot)$ denotes the DINOV2 feature extractor, $\mathcal{M}_{\mathrm{obj}}(v)$ is the soft object mask at viewpoint $v$ obtained from SAM2, and $\mathcal{M}_{\mathrm{grip}}$ is the fixed soft gripper mask. The first term enforces feature alignment, while the second penalizes overlap between the object and gripper masks. Unlike photometric losses employed in prior work [yen2021inerf], our loss explicitly accounts for occlusions. We note although DINOV2 is unreliable for global initialization, it is effective for local refinement.
@@ III-D View-Conditioned Imitation Learning
AV Trajectory Transfer. Trajectory transfer (TT) methods estimate the relative pose change of the object between the demonstration and the test instance, and then transfer the demonstration trajectory in a one-shot manner within the $\mathrm{SE}(3)$ manifold [doi:10.1126/scirobotics.adv7594]. Consequently, only a single demonstration optimal view is required. In our setting, demonstrations are recorded in the actor arm’s coordinate frame, while observations are captured from the observer arm’s camera. To resolve this mismatch, we extend trajectory transfer by applying the following sequence of frame transformations:
@@
AV Behavior Cloning. Behavior cloning (BC) trains a closed-loop policy $\pi$ to map observations to actions. Unlike TT, which exploits pose estimation models, BC requires several demonstration optimal views to generalize across diverse camera perspectives covering $v^{*}_{\mathrm{test}}$. Since optimizing over all viewpoints in $\mathcal{V}^{*}_{\mathrm{train}}$ is computationally expensive, we use Mast3R to estimate camera poses in the object frame and select the viewpoint closest to the center of all estimated poses as $v^{*}_{\mathrm{demo}}$, which we found to work well empirically. Moreover, unlike prior work [pmlr-v270-cheng25b, chuang2024active], which represents the end-effector pose in a static world frame and provides the camera pose explicitly as policy input, we express the actor arm’s end-effector pose directly in the camera frame:
@@ IV Evaluation
We instantiated our method on a real-world dual-arm ALOHA setup [zhao2023learning], mounted on a table and equipped with two calibrated RealSense D405 cameras. To evaluate our approach, we selected five diverse manipulation tasks involving self-occluding objects. In Pick Mug, the robot grasps a mug by its handle. In Hammer Nail, it uses a pre-grasped hammer to drive a nail. Open Drawer requires opening the second drawer of a multi-layer unit. In Retrieve Pack, the robot retrieves a package from a deep box. Finally, Insert Coin involves placing a coin into a storage container. All tasks require observing specific object parts or features for successful manipulation, with the key regions highlighted in red in Figure 3.
@@ IV-A Comparison with Static-Camera
To ensure fairness, we collect separate datasets for BC with and without active vision, rather than reusing the same dataset as was done in [chuang2024active]. Reusing would introduce visual inconsistencies, as the static camera would capture the observer arm moving across different poses, providing irrelevant information that could confuse the learned policy. In total, we collected 70 demonstrations for each of the five tasks, with and without active vision, resulting in $70\times 5\times 2=700$ demonstrations. All demonstrations were recorded using an iPhone-based teleoperation system built with ARKit.
We implement BC with ACT [zhao2023learning], using DINOv2 as the vision backbone and absolute Cartesian action representation. For BC with AV, we remove any frames in the dataset where the actor arm’s gripper is invisible in the camera view, as these introduce ambiguities. This also ensures that the trained policy outputs a pose where the actor’s gripper is likely to be visible in the camera view at the first timestep.
@@ IV-B Representing Actions in the Camera Frame
We evaluate the effect of representing actions in the camera frame for BC with AV. We compare against prior approaches [chuang2024active, xiong2025vision], where actions are represented in the fixed actor robot frame and the camera pose is provided as an input. Results in Table II show that representing actions directly in the camera frame improves both generalization and task success. We hypothesize that providing the camera pose as input is less effective because object poses span a wide distribution while camera views remain relatively similar, making it difficult for the policy to infer the robot’s pose in the static frame. On the other hand, representing actions directly in the camera frame provides a consistent reference.

### v2

@@ I Introduction
Current imitation learning methods for robotic manipulation primarily rely on static cameras [6, 10], egocentric wrist-mounted cameras [7, 9], or a combination of both [38, 30]. Static cameras are typically placed in a task-agnostic manner, and once a policy is trained, it often must be deployed in the same setup. Wrist-mounted cameras offer greater flexibility but suffer from limited global awareness and a restricted field of view. Combining both viewpoints can partially mitigate these issues but often introduces redundant observations that may distract the model—especially in low-data regimes [8, 32]. These limitations are further amplified by occlusions, highlighting the need for active vision, where the camera can move dynamically to provide the best viewpoint for the task.
Realizing active vision requires both movable camera hardware and sophisticated active vision strategies. Recent studies [8, 32, 5] achieve this by developing a dedicated active-vision arm for perception, with its policy learned through teleoperation and behavior cloning. However, this approach has two significant limitations. First, the active vision arm is restricted to a fixed observer role, limiting its reachable viewpoints and preventing its use as regular manipulator. Second, the active vision strategy demands extensive human demonstrations, which, although collectable via a VR-based teleoperation system, imposes additional burden on operators.
In this paper, we present ObAct (short for Observer–Actor), a novel framework for active vision in imitation learning, where an observer robot computes optimal visual observations that guide the actor robot’s actions. As illustrated in Figure LABEL:fig:intro with a mug-handle pickup task, ObAct positions the observer arm at test-time optimal viewpoints, providing visual observations for the actor arm during manipulation. Unlike prior approaches, ObAct dynamically assigns observer and actor roles based on the scene configuration, without requiring a separately trained active vision strategy for each arm. To enable this, we construct a test-time 3D Gaussian Splatting (3DGS) representation [16] from sparse-view images and optimize the camera pose to obtain observations that are similar to the demonstration observations, while minimizing occlusions. We further extend existing imitation learning methods—trajectory transfer [10] and behavior cloning [38]—to leverage this view-conditioned framework. Experiments show that ObAct substantially improves success rates over static-camera baselines: trajectory transfer increases by 145% without occlusion and 233% with occlusion, while behavior cloning increases by 75% and 143%, respectively.
@@ II Related Work
Active Vision for Robotic Manipulation. Active vision (AV) [1] refers to actively moving the camera to maximize perceptual quality for downstream tasks, and has been widely applied in robotics for object recognition [31], pose estimation [31, 2], and 3D reconstruction [15, 18]. It has also been extensively studied in robotic grasping, where unobserved views are sampled and evaluated using depth images based on information gain [3], grasp affordances [37], graspness inconsistency [20], and velocity fields [26], yielding improved performance in occluded and cluttered environments. In contrast, our method renders and evaluates unobserved viewpoints directly in the RGB domain, the prevalent modality in modern imitation learning systems [6, 38], and extends beyond grasping to support a broader range of manipulation tasks. Our setup is most closely related to [29], where an agent first observes the environment and then interacts with it by jointly training a next-best-view camera policy and a next-best-pose gripper policy using few-shot reinforcement learning. However, their approach assumes a freely moving camera and is validated only in simulation, whereas we demonstrate a real-world dual-arm system that achieves active vision without requiring a separate policy. In concurrent work, [33] also avoid training by leveraging 3DGS to build a scene model from dense scans and applying Bayesian optimization to determine an initial robot pose for task execution on a mobile platform. However, our focus is on sparse-view inputs in a dual-arm setting. [17] capture demonstrations with a 360° camera and jointly train an eyeball gaze controller and arm policies using a BC–RL loop, but their 2-DoF gaze mechanism is less capable of handling occlusions compared to our 6-DoF observer arm.
@@
View-conditioned Imitation Learning. For manipulation systems to generalize, visuomotor policies must remain robust to discrepancies in observational viewpoints—particularly unseen camera poses—which is especially critical in active vision settings where camera positions are dynamic rather than static. Recent works address this by training end-to-end behavior cloning policies that jointly control both the active camera and the manipulation arms using human demonstrations [32, 5, 8]. While effective, these approaches increase data requirements and model complexity. In contrast, our approach trains policies over a compact distribution of viewpoints and explicitly optimizes the test-time viewpoint relative to demonstrations. Another class of methods builds explicit 3D scene representations, such as point clouds or voxels [27, 36, 13], or renders canonical 2D views from these structures [14]. These approaches improve robustness to viewpoint variation but require precise extrinsic calibration and incur high computational costs. Data augmentation offers a complementary direction: [28] generate multi-view datasets from single-view demonstrations using a pre-trained zero-shot novel-view synthesis model [25], while [34, 24] edit reconstructed 3D scenes built from dense scans to create synthetic training views. Beyond end-to-end training, some works transfer a single demonstration trajectory at test time by estimating relative object pose changes between demonstration and execution views, either via pose estimation [10] or visual servoing [30]. However, such methods remain highly sensitive to occlusions. In this paper, we extend behavior cloning and trajectory transfer within our observer–actor framework, demonstrating consistent performance gains in both non-occluded and occluded scenarios.
@@ III-B Sparse-view Gaussian Splatting
Role Assignment. Roles are assigned to each arm based on how closely their captured views match $v^{*}_{\mathrm{demo}}$. Instead of performing explicit relative pose estimation, we use the robust dense feature matcher RoMa [11], using the number of confident matches on the segmented object as a proxy. The arm with more aggregated matches is designated as the observer, indicating it is closer to the demonstration view, while the other arm becomes the actor.
Three-View GS. We reconstruct the scene using the three images captured by the observer, via InstantSplat [12], a sparse-view method that leverages geometric priors. It employs Mast3R [19], a large-scale 3D geometric model, to estimate camera poses from RGB images and perform joint optimization. These optimized poses are more accurate than those from our low-cost robot arms, resulting in higher-quality reconstructions. We use only three images to train the 3DGS model, as the additional views provide little extra information for rendering $v^{*}_{\mathrm{test}}$, and training is considerably faster. We demonstrate that three views offer the best trade-off between performance and computation time (Section IV-C).
@@ III-C View Optimization
Candidate View Sampling. To initialize the optimization, we sample a hemisphere of candidate viewpoints around the task object’s center, following [3]. The object center is estimated from its 3D bounding box, obtained by lifting 2D masks from GroundedSAM [23] into 3D using aligned depth maps. After filtering out kinematically infeasible poses, we obtain a candidate set $\mathcal{V}_{\mathrm{candidate}}\subseteq\mathcal{V}$. Owing to the efficiency of 3DGS, all candidate views can be rendered at $\sim$250 FPS.
Optimal View Initialization. After rendering a set of virtual RGB images, we select the candidate view with the highest score as the initialization. The scoring function, identical to that used for role assignment, is based on the number of confidence-weighted feature matches between images from $\mathcal{V}_{\mathrm{candidate}}$ and $v^{*}_{\mathrm{demo}}$. This criterion implicitly accounts for occlusion, as occluded regions naturally produce fewer and less reliable matches. In practice, we found that this matching-based strategy outperforms an alternative based on DINOV2 [21] cosine distance for measuring view similarity. Even when combined with object segmentation, DINOV2 struggles to discriminate between viewpoints—an observation consistent with [4]. To reduce computational cost, we employ Tiny RoMa [11], a lightweight variant of RoMa with faster runtime.
Differentiable Rendering. After obtaining an initialization, we refine the viewpoint $v$ using differentiable rendering while explicitly accounting for gripper-induced occlusions. This is particularly important in our setting, where the observer arm’s gripper consistently remains in the field of view and may block the object, as shown in Figure 3. To handle this, we use SAM2 [22] for segmentation of the rendered image, guided by the demonstration image’s mask as a prompt. This is effective because the initialized and demonstration viewpoints are already closely aligned. To mitigate errors in estimating the object center, we first re-center the object’s mask in the image plane. We then refine $v$ by optimizing a loss that aligns the image rendered from the 3DGS model (denoted $\mathcal{G}$) with the segmented demonstration image $I^{*}_{\mathrm{demo}}$, while penalizing occlusions. Let $I(v;\mathcal{G})$ denotes the rendered image. The loss $\mathcal{L}$ as a particular view $v$ is defined as:
@@
where ${\phi}(\cdot)$ denotes the DINOV2 feature extractor, $\mathcal{M}_{\mathrm{obj}}(v)$ is the soft object mask at viewpoint $v$ obtained from SAM2, and $\mathcal{M}_{\mathrm{grip}}$ is the fixed soft gripper mask. The first term enforces feature alignment, while the second penalizes overlap between the object and gripper masks. Unlike photometric losses employed in prior work [35], our loss explicitly accounts for occlusions. We note although DINOV2 is unreliable for global initialization, it is effective for local refinement.
@@ III-D View-Conditioned Imitation Learning
AV Trajectory Transfer. Trajectory transfer (TT) methods estimate the relative pose change of the object between the demonstration and the test instance, and then transfer the demonstration trajectory in a one-shot manner within the $\mathrm{SE}(3)$ manifold [10]. Consequently, only a single demonstration optimal view is required. In our setting, demonstrations are recorded in the actor arm’s coordinate frame, while observations are captured from the observer arm’s camera. To resolve this mismatch, we extend trajectory transfer by applying the following sequence of frame transformations:
@@
AV Behavior Cloning. Behavior cloning (BC) trains a closed-loop policy $\pi$ to map observations to actions. Unlike TT, which exploits pose estimation models, BC requires several demonstration optimal views to generalize across diverse camera perspectives covering $v^{*}_{\mathrm{test}}$. Since optimizing over all viewpoints in $\mathcal{V}^{*}_{\mathrm{train}}$ is computationally expensive, we use Mast3R to estimate camera poses in the object frame and select the viewpoint closest to the center of all estimated poses as $v^{*}_{\mathrm{demo}}$, which we found to work well empirically. Moreover, unlike prior work [5, 8], which represents the end-effector pose in a static world frame and provides the camera pose explicitly as policy input, we express the actor arm’s end-effector pose directly in the camera frame:
@@ IV Evaluation
We instantiated our method on a real-world dual-arm ALOHA setup [38], mounted on a table and equipped with two calibrated RealSense D405 cameras. To evaluate our approach, we selected five diverse manipulation tasks involving self-occluding objects. In Pick Mug, the robot grasps a mug by its handle. In Hammer Nail, it uses a pre-grasped hammer to drive a nail. Open Drawer requires opening the second drawer of a multi-layer unit. In Retrieve Pack, the robot retrieves a package from a deep box. Finally, Insert Coin involves placing a coin into a storage container. All tasks require observing specific object parts or features for successful manipulation, with the key regions highlighted in red in Figure 3.
@@ IV-A Comparison with Static-Camera
To ensure fairness, we collect separate datasets for BC with and without active vision, rather than reusing the same dataset as was done in [8]. Reusing would introduce visual inconsistencies, as the static camera would capture the observer arm moving across different poses, providing irrelevant information that could confuse the learned policy. In total, we collected 70 demonstrations for each of the five tasks, with and without active vision, resulting in $70\times 5\times 2=700$ demonstrations. All demonstrations were recorded using an iPhone-based teleoperation system built with ARKit.
We implement BC with ACT [38], using DINOv2 as the vision backbone and absolute Cartesian action representation. For BC with AV, we remove any frames in the dataset where the actor arm’s gripper is invisible in the camera view, as these introduce ambiguities. This also ensures that the trained policy outputs a pose where the actor’s gripper is likely to be visible in the camera view at the first timestep.
@@ IV-B Representing Actions in the Camera Frame
We evaluate the effect of representing actions in the camera frame for BC with AV. We compare against prior approaches [8, 32], where actions are represented in the fixed actor robot frame and the camera pose is provided as an input. Results in Table II show that representing actions directly in the camera frame improves both generalization and task success. We hypothesize that providing the camera pose as input is less effective because object poses span a wide distribution while camera views remain relatively similar, making it difficult for the policy to infer the robot’s pose in the static frame. On the other hand, representing actions directly in the camera frame provides a consistent reference.